Advertisements
Advertisements
प्रश्न
In an entrance test that is graded on the basis of two examinations, the probability of a randomly chosen student passing the first examination is 0.8 and the probability of passing the second examination is 0.7. The probability of passing at least one of them is 0.95. What is the probability of passing both?
Advertisements
उत्तर
Let A and B be the events of passing first and second examinations, respectively.
P(A) = 0.8, P(B) = 0.7
Probability of passing at least one examination
= 1 – P(A’ ∩ B’) = 0.95
⇒ P(A’ ∩ B’) = 1 – 0.95
= 0.05
But A’ ∩ B’ = (A ∪ B)’ ... (By Demorgan’s Law)
∴ P(A’ ∩ B’) = P(A ∪ B)’ = 1 – P(A ∪ B)
= 0.05
∴ P(A ∪ B) = 1 – 0.05
= 0.95
Now P(A ∪ B) = P(A) + P(B) – P(A ∩ B)
0.95 = 0.8 + 0.7 – P(A ∩ B)
P(A ∩ B) = 1.5 – 0.95
= 0.55
Thus, the probability of passing both the examinations = 0.55
APPEARS IN
संबंधित प्रश्न
A letter is chosen at random from the word ‘ASSASSINATION’. Find the probability that letter is
- a vowel
- an consonant
A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16. Determine P (not B)
A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16. Determine P(A or B).
In a class of 60 students, 30 opted for NCC, 32 opted for NSS and 24 opted for both NCC and NSS. If one of these students is selected at random, find the probability that
- The student opted for NCC or NSS.
- The student has opted neither NCC nor NSS.
- The student has opted NSS but not NCC.
Three letters are dictated to three persons and an envelope is addressed to each of them, the letters are inserted into the envelopes at random so that each envelope contains exactly one letter. Find the probability that at least one letter is in its proper envelope.
A dice is thrown. Find the probability of getting a prime number
A dice is thrown. Find the probability of getting a multiple of 2 or 3.
In a simultaneous throw of a pair of dice, find the probability of getting an even number on first
In a simultaneous throw of a pair of dice, find the probability of getting neither 9 nor 11 as the sum of the numbers on the faces
In a simultaneous throw of a pair of dice, find the probability of getting odd number on the first and 6 on the second
In a simultaneous throw of a pair of dice, find the probability of getting a total of 9 or 11
In a simultaneous throw of a pair of dice, find the probability of getting a total greater than 8.
In a single throw of three dice, find the probability of getting a total of 17 or 18.
Three coins are tossed together. Find the probability of getting at least two heads
Two dice are thrown. Find the odds in favour of getting the sum 4.
Two dice are thrown. Find the odds in favour of getting the sum 5.
A box contains 10 red marbles, 20 blue marbles and 30 green marbles. 5 marbles are drawn at random. From the box, what is the probability that all are blue?
A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is white .
A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is red or even numbered.
Fill in the blank in the table:
| P (A) | P (B) | P (A ∩ B) | P(A∪ B) |
| \[\frac{1}{3}\] | \[\frac{1}{5}\] | \[\frac{1}{15}\] | ...... |
Fill in the blank in the table:
| P (A) | P (B) | P (A ∩ B) | P(A∪ B) |
| 0.5 | 0.35 | ..... | 0.7 |
If A and B are two events associated with a random experiment such that
P(A) = 0.5, P(B) = 0.3 and P (A ∩ B) = 0.2, find P (A ∪ B).
If A and B are two events associated with a random experiment such that
P (A ∪ B) = 0.8, P (A ∩ B) = 0.3 and P \[(\bar{A} )\]= 0.5, find P(B).
There are three events A, B, C one of which must and only one can happen, the odds are 8 to 3 against A, 5 to 2 against B, find the odds against C
A card is drawn at random from a well-shuffled deck of 52 cards. Find the probability of its being a spade or a king.
In a single throw of two dice, find the probability that neither a doublet nor a total of 9 will appear.
100 students appeared for two examination, 60 passed the first, 50 passed the second and 30 passed both. Find the probability that a student selected at random has passed at least one examination.
The probability that a leap year will have 53 Fridays or 53 Saturdays is
If the probability of A to fail in an examination is \[\frac{1}{5}\] and that of B is \[\frac{3}{10}\] . Then, the probability that either A or B fails is
A box contains 10 good articles and 6 defective articles. One item is drawn at random. The probability that it is either good or has a defect, is
A bag contains 5 black balls, 4 white balls and 3 red balls. If a ball is selected randomwise, the probability that it is black or red ball is
A box contains 10 good articles and 6 with defects. One item is drawn at random. The probability that it is either good or has a defect is
Three numbers are chosen from 1 to 20. The probability that they are not consecutive is
